1. Calculate the pH and pOH of a 0.036 M solution of HBr.
Answer #1

HBr is a strong acid.

[H+] = [HBr] = 0.036 M

pH = -log[H+] = -log(0.036) = -(-1.44) = 1.44

pOH = pKw - pH = 14.00 - 1.44 = 12.56